OpenStudy (anonymous):

What is the probability of getting tails each time if a coin is tossed 4 times?

OpenStudy (kidrah69):

hey :) ok so how many sides does a coin have?

OpenStudy (anonymous):

2

OpenStudy (rational):

whats the probability of getting a tail the first time ?

OpenStudy (anonymous):

$$ \Large P(TTTT) = P(T) \times P(T) \times P(T) \times P(T) $$